# DSAI: Unbiased and Interpretable Latent Feature Extraction for Data-Centric AI

### v1

@@ 1 Introduction
Figure 1: DSAI automates the extraction of latent features like syntax, domain, and clarity
@@
The DSAI pipeline operates in five stages: #1 Perspective Generation identifies data-driven perspectives from a small subset of data. #2 Perspective-Value Matching assigns values to individual data points by evaluating them against these perspectives. #3 Clustering groups values with shared characteristics to reduce redundancy. #4 Verbalization converts extracted features into a compact criterion form. #5 Prominence-based Selection determines which features to use based on a prominence intensity metric that quantifies the discriminative power of each extracted feature. Throughout the process, LLMs remain task-agnostic to minimize bias, ensuring that their outputs are grounded in data rather than domain-specific assumptions.
To validate our framework, we demonstrate its effectiveness on datasets specifically curated for our experiments, which include research titles Wang et al. (2018) and slogans Jin et al. (2023). We then showcase DSAI’s practical value using real-world datasets: news headlines with CTR Wu et al. (2020), spam detection Kim (2016), and Reddit comments with community engagement metrics Magnan (2019). These experiments highlight DSAI’s ability to provide actionable insights for industries focused on content optimization, user engagement, and moderation.
Our main contributions are as follows: (1) Minimizing bias by ensuring LLMs identify latent characteristics within large datasets; (2) Introducing a quantitative metric for feature prominence scoring, which serves as a proxy for measuring discriminative power of each feature; (3) Improving interpretability through feature-to-source traceability. (4) Enabling thorough examination of large datasets while minimizing human labor through systematic guidance of LLMs;
@@ 3 Challenges in LLM-Driven Data Analysis
This section demonstrates that LLMs rely more on pre-trained knowledge than data-specific features, showing a tendency to depend on pre-trained knowledge when input data is absent or labels are flipped, and generating vague features when task-specific context is removed.
3.1 Setting
Dataset Annotation and Sampling
Our experimental task aims to identify the characteristics of high-quality text as defined by expert criteria Nair and Gibbert (2016); Tullu (2019); Kohli et al. (2007); Padrakali and Chitra Chellam (2017). We use two distinct datasets: (1) commercial slogans and (2) academic research titles, representing different text genres (Appendix A).
To construct the high-quality sample set, we used LLMs to evaluate each entry in the 3,000-entry dataset according to expert criteria. Binary scores were assigned for each criterion, and final ratings were aggregated. The top 600 samples formed the analysis group ("positive"), and the bottom 600 formed the control group ("negative").222”Positive”/”negative” indicates the presence/absence of specific features rather than overall quality. See Appendix C for detailed annotation procedures.
Model
We used GPT-4o OpenAI (2024) for its advanced capabilities in nuanced text understanding and annotation tasks Tan et al. (2024).
3.2 Data Grounding Issues
We explore directly extracting latent features using LLMs with two different types of input data:
  • PosData: Uses only positive (top-scored) samples to identify characteristics of high-quality samples.
  • MixedData: Uses both positive (top-scored) and negative (bottom-scored) samples to capture a broader range of features.
Both approaches have generated features that substantially overlap with expert-defined criteria (Table 1). However, since our datasets come from popular domains, it is unclear whether the LLM-generated features are truly derived from the provided data or from the model’s pre-existing domain knowledge. To test this, we pose several key questions and conduct experiments to evaluate them.
@@ 3.2 Data Grounding Issues
We tested whether the LLM uses input data to adapt its feature generation. To evaluate this, we flip the labels and observe if the generated features change accordingly:
  • FlippedPosData: Negative examples are labeled as positive.
  • FlippedMixedData: Positive and negative labels are reversed.
As shown in Table 1, features generated from PosData and FlippedPosData exhibited minimal changes, with significant overlap also observed between MixedData and FlippedMixedData. This indicates that LLMs rely more on pre-existing knowledge than on input data for feature generation.
@@
To determine if features are primarily sourced from internal knowledge, we generated features without any input data:
  • NoData: The LLM generates features solely based on its pre-trained knowledge.
Table 1 shows that NoData achieved high recall on expert-defined criteria and significant overlap with PosData, MixedData, and flipped datasets. These findings confirm that pre-existing knowledge, rather than data-specific information, drives LLM-generated features.
@@
We computed the discriminative power (DP) of features to assess whether they reflect the latent characteristics of positive data (details in Section 5.2). Features with DP values $\leq 0.5$ fail to differentiate between positive and negative traits, while values $>0.6$ indicate strong discrimination.
As shown in Figure 2(a), several features from PosData and MixedData exhibited low DP scores, some below 0.5. These results highlight the limitations of LLMs in generating features that truly represent positive data’s latent traits.
@@
To test if removing task-specific context reduces bias and improves data grounding, we used:
  • NoContext: The LLM receives only unlabeled data groups and is unaware of the goal of identifying high-quality text.
In this setting, the LLM defaulted to generating vague and generic descriptors (Appendix D). This suggests that task context removal alone does not improve data grounding and highlights the need for structured guidance during feature extraction.
@@ 4 Data Scientist AI
DSAI is a 5-stage framework for automated latent feature generation, leveraging LLM capabilities with structured processes (Figure 3). Below, we outline the DSAI pipeline and its five stages.
@@
LLMs can extract diverse attributes from text Peng et al. (2023). Using a subset of positive and negative samples, this stage generates distinct attributes while minimizing reliance on domain bias by excluding task-specific context. Each perspective includes a name, evaluation criterion, analysis process, and example (Figure 3 (1)). This structure provides a consistent framework for dataset interpretation. After generating perspective candidates, duplicates are removed.
@@
This stage assigns values to each data point guided by the generated perspectives on the previous stage.
@@
This stage groups similar values within each perspective to create interpretable feature sets with reduced redundancy. Based on LLM clustering literature Wang et al. (2023); Wan et al. (2024); Lam et al. (2024), the process consists of two steps: generating representative cluster labels, and assigning values to appropriate clusters (Figure 3 (3)).
@@
This stage transforms perspective-label pairs into verbal criteria. For each pair, we compute $P(\textrm{positive}|D_{p,l})$, the proportion of positive examples in dataset $D_{p,l}$ corresponding to the perspective-label pair $(p,l)$. Using this, we calculate a directional score $12\times P(\textrm{positive}|D_{p,l})-$ 1, which determines how the pair should be verbalized. Pairs with positive directional scores (>0) are directly verbalized as features describing positive data. Pairs with negative directional scores (<0) are transformed into "avoid" statements, which indirectly characterize positive data by specifying features to avoid. [^3] For instance, if (clarity, low) receives a negative directional score, it is verbalized as ”Avoid sentences with low clarity.” This dual transformation approach ensures coverage of both desired and undesired traits.
@@
Finally, DSAI employs prominence intensity as the feature selection metric, defined as the absolute value of directional score $\|2*P(\textrm{positive}|D_{p,l})-1\|$. This metric indicates how decisively a perspective-label pair characterizes either positive or negative samples–the closer to 1, the stronger the association. Unlike our baseline approaches (Section 3.2) that merely output a list of features without any relative importance measures, our prominence intensity metric provides a quantitative way to assess and compare the discriminative power of each feature. Users can set a prominence intensity threshold to match their specific needs, balancing between stronger predictive patterns (higher threshold) and broader coverage of the dataset’s latent characteristics (lower threshold).
Traceability and Transparency
The feature generation process maintains the link of each feature to its perspective-label combination, the original data samples that contributed to its creation, and the quantitative prominence score that justified its selection. Such traceability enables users to verify the legitimacy of extracted features by examining their source data and the complete derivation process.
@@ 5.1 Recall of Expert-Defined Criteria
We evaluate DSAI’s ability to reproduce expert-defined criteria by applying it to slogan and research title datasets annotated as described in Section 3.1.
@@
We generated criteria through our pipeline and retained those with $D_{p,l}>6$ and positive prominence intensity scores. This yielded 235 criteria for slogans and 198 criteria for research titles.
@@
Our methodology showed strong performance in reproducing expert-defined criteria, even at high prominence intensity thresholds. All 9 expert criteria for slogans were captured at a threshold of 0.348, and 83% recall (10/12) was achieved for research titles at a threshold as high as 0.692 (Table 1). While PosData and MixedData also showed decent coverage, their results relied on LLM’s pre-existing knowledge as discussed in Section 3.2. In contrast, our approach, by design, minimizes such potential bias by withholding task-specific context, while still achieving comparable or better recall rates.
@@
The adjustable prominence intensity thresholds allow users to tailor their analyses by balancing discriminative power and coverage. Additional analyses on recall dynamics across various thresholds are provided in Appendix H.
@@ 5.2 Discriminative Power (DP)
To validate our pipeline’s effectiveness, we examined whether criteria with higher prominence intensity scores indeed showed stronger discriminative power in practice. We divided DSAI-generated criteria into five buckets based on prominence intensity scores to cover the full range. From each bucket, 10 features were sampled and used to annotate datasets for feature conformity ("conform"/"not-conform") where applicable. Non-applicable cases444e.g., features referencing business models in unrelated data. were excluded from calculations.
We define DP score as follows:
$DP\text{ }Score=\left\{\begin{aligned} &P(\textrm{positive}|\textrm{conform})\quad\text{if }\\ &\quad\quad\quad\quad\quad\textrm{directional\_score}>0,\\ &P(\textrm{negative}|\textrm{notConform})\quad\text{elif }\\ &\quad\quad\quad\quad\quad\textrm{directional\_score}<0.\end{aligned}\right.$
DP scores above 0.5 indicate effective discrimination of positive examples, while scores below 0.5 suggest potential errors in the LLM-based pipeline (either in direct or "avoid" form). Criteria scoring between 0.5 and 0.6 were deemed too generic to provide meaningful discrimination.
As shown in Figure 2(b), criteria with higher prominence scores demonstrated stronger discriminative power, with all DSAI-generated criteria scoring above 0.5. This validates not only our pipeline’s reliability but also the effectiveness of the prominence intensity metric in identifying meaningful discriminative criteria. Through this metric, users can quantitatively assess which features are more reliable indicators of positive samples, rather than treating all extracted features as equally important.
5.3 Reliability of Pipeline Operations
Due to the inherent uncertainty in LLM outputs, we evaluated the reliability of pipeline stages that rely on LLM operations by having the LLM verify its own mappings between datapoints and features. We assessed three LLM-dependent stages by asking the LLM to confirm the reliability of its outputs. [^5] Stage #1 (Perspective Generation) was not evaluated as perspective generation allows for more diverse outputs. Stage #5 (Prominence-based Selection) was excluded as the results are deterministic. Our verification process showed high consistency rates of >98%,  94%, and 98% for stages #2, #3, and #4 respectively (details in Appendix J).
6 Real-World Application with Quantitative Datasets
We applied DSAI to three real-world user feedback datasets critical for business insights Luo et al. (2022): (1) MIND Wu et al. (2020), analyzing engagement features in news headlines with high CTR as the positive group; (2) spam detection Kim (2016), identifying patterns in spam messages as the positive group; and (3) Reddit Magnan (2019), exploring interaction-promoting linguistic features in highly upvoted comments as the positive group.
Prominence Distribution and Samples
The datasets show distinct patterns, highlighting the need for dataset-specific thresholds to balance discriminative power, recall, and domain priorities (e.g., minimizing spam errors or maximizing engagement). In addition to general characteristics, DSAI-generated features notably capture more specific and fine-grained nuances across various subsets of a given dataset, often overlooked in LLM-based analyses. Detailed analysis is provided in Appendix L.
@@ 6 Real-World Application with Quantitative Datasets
Our methodology provides latent characteristics of data in linguistic form, enabling diverse applications such as style transfer, annotation guideline generation, and classification tasks. To demonstrate the practical utility of DSAI-generated criteria, we conducted a toy spam classification experiment using 20 spam and 20 ham samples. Using the five criteria with the highest prominence intensity scores led to effective classification performance, while using the five criteria with the lowest prominence resulted in poor performance (See Figure 4). This illustrates DSAI’s potential to establish interpretability along with feature credibility through prominence scores, making it suitable for automatic application in various downstream tasks including classification.

### v2

@@ 1 Introduction
The DSAI pipeline operates in five stages: #1 Perspective Generation identifies data-driven perspectives from a small subset of data. #2 Perspective-Value Matching assigns values to individual data points by evaluating them against these perspectives. #3 Clustering groups values with shared characteristics to reduce redundancy. #4 Verbalization converts extracted features into a compact criterion form. #5 Prominence-based Selection determines which features to use based on a prominence intensity metric that quantifies the discriminative power of each extracted feature.
Throughout these stages, the LLM remains task-agnostic – we do not reveal the specific domain or the “correct answer” context during feature generation. This design minimizes bias and ensures that the identified features are grounded in the data rather than the model’s background knowledge.
We validate our framework on datasets curated for our experiments, including a research titles dataset Wang et al. (2018) and an advertising slogans dataset Jin et al. (2023), each with expert-defined criteria that serve as ground truth for evaluation. We then demonstrate DSAI’s practical value on three real-world datasets: news headlines with click-through rate (CTR) labels Wu et al. (2020), a spam detection dataset Kim (2016), and Reddit comments with community engagement metrics Magnan (2019).
Our main contributions are as follows:
  • Minimizing Bias: We ensure that LLMs focus on latent characteristics present in the data, thus mitigating the tendency to rely on domain-specific prior knowledge (addressing the data grounding issue).
  • Prominence Metric: We introduce a quantitative metric for feature prominence, which serves as a proxy for the discriminative power of each extracted feature.
  • Interpretability: We improve interpretability through feature-to-source traceability, allowing users to trace each extracted feature back to the data points that support it.
  • Efficiency: We enable thorough examination of large datasets with minimal human labor by systematically guiding LLMs through the analysis process.
@@ 3 Challenges in LLM-Driven Data Analysis
In this section, we examine the behavior of a state-of-the-art LLM when tasked with directly generating features from data. These exploratory experiments reveal the data-grounding challenges that motivate our DSAI approach.
3.1 Setting
Dataset Annotation and Sampling
We use two expert-annotated text datasets where the “ground truth” latent features are known (defined by domain experts). The first dataset consists of research paper titles and the second contains advertising slogans. In each dataset, experts Nair and Gibbert (2016); Tullu (2019); Kohli et al. (2007); Padrakali and Chitra Chellam (2017) have outlined a set of specific criteria or features that good examples should exhibit. We annotated each sample for the presence or absence of each criterion. This yields a detailed profile of which features are present in each data point. We then designated positive examples as those that satisfy many of the expert criteria (top-scoring "high-quality” samples) and negative examples as those that violate or lack many of the criteria (bottom-scoring "low-quality” samples). This dataset construction emphasizes concrete feature differences rather than a subjective quality judgment.
Model
All experiments in this section use GPT-4o OpenAI (2024) as the LLM, which is shown to have strong capabilities in understanding nuanced text and performing annotation-like tasks Tan et al. (2024). The model was prompted in a zero-shot manner to generate dataset features under various conditions.
3.2 Data Grounding Issues
@@
Using the above datasets, we explored straightforward ways of prompting the LLM to extract latent features, and evaluated whether the LLM’s outputs were truly grounded in the input data. We tried two input configurations for the prompt:
  • PosData: Provide the LLM with only positive examples and ask it to identify characteristics common to these examples.
  • MixedData: Provide the LLM with both positive and negative examples, and ask it to find distinguishing features.
In both cases, we formatted the prompt to list a set of representative samples and requested the model to output a list of key features or criteria describing the positive group. We found that both PosData and MixedData prompts yielded feature lists that substantially overlapped with the expert-defined criteria (Table 1). On the surface, this suggests the model can produce reasonable-sounding features. However, because our datasets come from well-known domains, we must question whether the LLM actually derived these features from the given data, or if it simply regurgitated its own prior knowledge about the domain. To investigate this, we designed four experiments around the following questions:
@@
We tested whether the LLM truly uses the input examples to adapt its feature generation. If the model is grounding its output in the provided data, then changing the data labels should lead to corresponding changes in the generated features. To check this, we flipped the class labels in the input and observed the effect on the output features:
  • FlippedPosData: We took the negative examples but misled the LLM by labeling them as if they were positive ("high-quality") examples.
  • FlippedMixedData: We presented the LLM with the same set of positive and negative examples as MixedData, but swapped their labels (positives labeled as “low-quality” and negatives labeled as “high-quality”).
If the model adapted to the data, FlippedPosData should produce features for low-quality texts, and FlippedMixedData should invert the original MixedData features. Instead, we observed the opposite. As shown in Table 1, FlippedPosData produced features nearly identical to PosData, and FlippedMixedData closely matched MixedData. This suggests that flipping labels had minimal impact, indicating the model relied on prior notions of “high-quality” text rather than adapting to the input data.
@@
The above result raises the question: would the LLM have generated a similar list of features even if we gave it no data at all? To test this, we prompted the LLM to list features of high-quality text without providing any example data. Here, the model must rely solely on its internal knowledge.
  • NoData: The LLM is asked to imagine or recall what characteristics define good content, without seeing any specific dataset samples.
Remarkably, NoData features showed high recall of expert-defined criteria, strongly overlapping with PosData and MixedData (Table 1). Even without input data, the model reproduced nearly all expert criteria—fully for slogans and almost entirely for titles. This suggests the LLM relied on prior knowledge rather than capturing from the provided data, highlighting both its strong knowledge base and limited data-driven adaptation. While this demonstrates the model’s impressive knowledge base, it also underscores the lack of true data grounding in these direct generation approaches.
@@
We next examined the quality of the features generated by the LLM in terms of how well they actually characterize the positive class in the data. Just because a feature sounds like a good guideline doesn’t guarantee that it differentiates our positive and negative examples. For instance, the model might output “Use simple language” as a feature of good text, but if both our positive and negative examples equally exhibit (or fail to exhibit) this trait, then that feature isn’t really capturing what makes the positive group unique in our dataset.
[…]
Using this metric, we evaluated the features generated under the PosData and MixedData prompts. We found that several of those features had low DP scores, some even below 0.5 (Figure 1(a)). This means the model sometimes proposed features that were more prevalent in the “bad” examples than the “good” ones. These results highlight a limitation of directly using an LLM for feature generation: many of the features it outputs, even if they sound plausible, do not truly reflect the distinguishing characteristics of the positive data.
@@
One hypothesis to improve grounding was that the LLM’s knowledge might be overly triggered by the context of the task in the prompt. To test if removing task-specific context reduces bias and improves data grounding, we used:
  • NoContext: The LLM is given two unlabeled sets of texts but is not told which set is “high-quality” or “low-quality”, nor even that the goal is to identify high-quality text features. Essentially, we ask the model to compare two groups of texts without naming the task.
In this setting, the LLM defaulted to generating vague and generic descriptors (Appendix LABEL:appendix:baseline6). This suggests that task context removal alone does not improve data grounding and highlights the need for structured guidance during feature extraction.
@@ 4 Data Scientist AI
DSAI is a five-stage framework for automated latent feature generation that leverages LLM capabilities within a structured process. The pipeline, illustrated in Figure 2, is designed to overcome the shortcomings identified in §3.2 by guiding the LLM through controlled steps.
@@
In this first stage, we prompt the LLM to generate a diverse set of perspectives – these are different angles or aspects under which the data can be described. The key idea is to break down the analysis into multiple facets. To achieve this, we feed the LLM a small subset of the dataset, including a few positive and negative examples. By keeping the task context hidden, we minimize domain bias while still providing concrete data for the model to analyze. We then ask the LLM to propose distinct perspectives that might explain differences in the data.
Each perspective generated by the LLM comes with a structured description: we instruct the model to give each perspective a name, a brief evaluation criterion, a suggested process for analyzing a text from that perspective, and an example from the provided data illustrating the perspective in action (Figure 2 (1)). By having the LLM produce this structured output, we ensure consistency in how it conceptualizes each aspect. After generation, we also apply a de-duplication step to remove redundant or overlapping perspectives. The result of Stage 1 is a list of candidate perspectives, which serve as the conceptual foundation for the subsequent analysis.
@@
In Stage 2, we systematically evaluate each data point against the perspectives identified in Stage 1. For every (perspective, data point) pair, the LLM is prompted to assign a value that describes the data point with respect to that perspective. After this stage, each data point is associated with a set of (perspective, value) pairs – one for each perspective – capturing how the LLM perceives that point on each aspect.
@@
To make the features more interpretable and reduce redundancy, we cluster similar values within each perspective. We employ the LLM for this clustering task, drawing on its semantic understanding. The process works in two sub-steps for each perspective: (1) Cluster Label Generation: The LLM examines all the values it assigned under a given perspective and proposes a smaller set of representative labels (cluster names). (2) Value Assignment: Next, the LLM assigns each raw value to one of the generated cluster labels.
The outcome is that for each perspective, we now have a handful of feature categories (the cluster labels), each representing a group of similar values. Each data point can thus be described in terms of these cluster labels.
@@
This stage transforms (perspective, label) pairs into verbal criteria. For each pair, we compute $P(\textrm{positive}|D_{p,l})$, the proportion of positive examples in dataset $D_{p,l}$ corresponding to the (perspective, label) pair $(p,l)$. Using this, we calculate a directional score $12\times P(\textrm{positive}|D_{p,l})-$ 1, which determines how the pair should be verbalized. Pairs with positive directional scores (>0) are directly verbalized as features describing positive data. Pairs with negative directional scores (<0) are transformed into "avoid" statements, which indirectly characterize positive data by specifying features to avoid. [^2] For instance, if (clarity, low) receives a negative directional score, it is verbalized as ”Avoid sentences with low clarity.” This dual transformation approach ensures coverage of both desired and undesired traits.
@@
Finally, DSAI employs prominence intensity as the feature selection metric, defined as the absolute value of directional score $\|2*P(\textrm{positive}|D_{p,l})-1\|$. Using prominence as a metric, we can now select the most impactful features by setting a prominence threshold (which can be adjusted by the user) and retain only features above that threshold. The key is that DSAI is not just outputting an unstructured list of features – it provides a way to prioritize them. This addresses the issue with the baseline LLM approach (§3.2), which gave no indication of which features were more important or reliable. By looking at the prominence scores, users can decide how many features to consider or where to draw the line between major and minor features.
After Stage 5, the final output of DSAI is a curated set of features, each in a clear natural-language form, typically accompanied by their prominence scores. These features are intended to be data-grounded and interpretable, providing insight into the data. In the next section, we evaluate how well this pipeline works in practice, especially in comparison to the direct LLM approach.
@@ 5 Validation of Methodology Using Expert-Driven Annotation Dataset
  Requirement Prominence Frequency
                         Top Requirements
                         Bottom Requirements
Ensure the use of inclusive language in the sentence. 0.0225 1,109
@@ 5.1 Recall of Expert-Defined Criteria
One way to gauge DSAI’s effectiveness is to see if it can rediscover the ground-truth criteria that domain experts have defined for these datasets. We applied the full DSAI pipeline to the slogans and research titles data annotated as described in §3.1.
@@
We generated criteria through our pipeline and retained those with $|D_{p,l}|>6$ and positive prominence intensity scores. This yielded 235 criteria for slogans and 198 criteria for research titles. Examples of the generated criteria are shown in Table 10.
@@
Our methodology showed strong performance in reproducing expert-defined criteria, even at high prominence intensity thresholds. All 9 expert criteria for slogans were captured at a threshold of 0.348, and 83% recall (10/12) was achieved for research titles at a threshold as high as 0.692 (Table 1). While PosData and MixedData also showed decent coverage, their results relied on LLM’s pre-existing knowledge as discussed in §3.2. In contrast, our approach, by design, minimizes such potential bias by withholding task-specific context, while still achieving comparable or better recall rates.
@@
The adjustable prominence intensity thresholds allow users to tailor their analyses by balancing discriminative power and coverage. We provide a detailed analysis of recall dynamics across different threshold values in Appendix D. In summary, we categorized expert-defined criteria, ranked their importance, and examined at which thresholds they were filtered out. More important criteria persisted at higher thresholds, while less critical ones were eliminated at lower thresholds.
@@ 5.2 Discriminative Power (DP)
Having shown that DSAI can reproduce known criteria, we next investigate whether the prominence score we assign to features actually correlates with real discriminative power. Intuitively, if our pipeline is working correctly, features with higher prominence intensity should be the ones that better distinguish positive from negative examples.
In our evaluation, we took all the DSAI-generated criteria and binned them into five buckets based on prominence intensity scores. From each bucket, we sampled 10 criteria for manual examination. For each sampled feature, we went back to the dataset and annotated instances to see if the feature was present or absent in those examples. We excluded non-applicable333e.g., feature about “use of sales terms” in examples unrelated to business models. cases from the calculations.
Using these annotations, we computed the DP score of each feature as mentioned earlier (§3.2(c)), and then examined the DP scores across the different prominence buckets.
$DP\text{ }Score=\left\{\begin{aligned} &P(\textrm{positive}|\textrm{feature-present})\quad\text{if }\\ &\quad\quad\quad\quad\quad\textrm{directional\_score}>0,\\ &P(\textrm{negative}|\textrm{feature-absent})\quad\text{elif }\\ &\quad\quad\quad\quad\quad\textrm{directional\_score}<0.\end{aligned}\right.$
The results confirmed our expectations: criteria with higher prominence scores generally showed higher DP. As illustrated in Figure 1(b), all of the DSAI-generated criteria achieved DP > 0.5. Lower-prominence features, in contrast, were sometimes borderline 0.5, reinforcing the idea that prominence is a good indicator of a feature’s reliability. In short, by using the prominence intensity metric, DSAI effectively filters and ranks features by their true discriminative power. This validation not only demonstrates the reliability of our pipeline’s outputs, but also highlights that the prominence metric can guide users to the most trustworthy features. Practically, this means one can focus on high-prominence features for critical decisions, knowing they have been quantitatively vetted to distinguish positive examples well.
5.3 Reliability of Pipeline Operations
DSAI’s multi-stage pipeline relies on the LLM’s output at several steps. While the previous sections show the end results are effective, we also wanted to ensure that each intermediate step was performed reliably by the LLM. To assess this, we had the LLM perform a self-check on its work for certain stages. Each verification was conducted in a separate session from the original task to ensure independence. We focused on the stages that have a well-defined objective where consistency can be measured:
  • Stage 2 (Value Matching): For each perspective and data point, after the LLM assigned a value, we asked the LLM to confirm whether that assignment was correct given the data point’s content.
  • Stage 3 (Clustering): After the LLM clustered values, we gave it each value along with the cluster label and asked if that assignment was appropriate.
  • Stage 4 (Verbalization): We asked the LLM to verify that each verbalized feature correctly described the intended cluster and perspective, and that the phrasing (direct or “avoid”) corresponded to the sign of the directional score.
The LLM’s self-audit verification process showed high consistency rates of >98%,  94%, and 98% for stages #2, #3, and #4 respectively. These results indicate that the pipeline’s internal operations are reliable; the LLM is largely consistent and does not contradict itself when asked to re-check its work.
6 Real-World Application with Quantitative Datasets
@@
Having validated DSAI on datasets with known ground truth criteria, we now apply it to several real-world datasets to demonstrate its practicality and versatility. We selected three real-world user feedback datasets critical for business insights Luo et al. (2022): (1) MIND Wu et al. (2020), analyzing engagement features in news headlines with high CTR as the positive group; (2) spam detection Kim (2016), identifying patterns in spam messages as the positive group; and (3) Reddit Magnan (2019), exploring interaction-promoting linguistic features in highly upvoted comments as the positive group. These datasets differ significantly in content and domain, which allows us to see how DSAI adapts to different domains without any domain-specific tuning.
Prominence Distribution and Sample Insights
Running DSAI on each dataset, we observed that the distribution of feature prominence scores differs by domain. This is expected: each domain has its own characteristics and noise levels, so the threshold for what constitutes a strongly discriminative feature will vary. For instance, in the spam dataset, it might be crucial to set a higher prominence threshold to avoid any features that could lead to false positives. In the news headline dataset, one might choose a threshold that balances identifying strong engagement drivers while not missing out on subtler but interesting patterns. The Reddit data might show a different spread, capturing nuances of informal language or humor that drive upvotes.
Importantly, DSAI captured not only general traits (like "uses urgent language" might be a spam trait common across many messages, or "mentions specific names/events" for news headlines) but also fine-grained nuances specific to subsets of each dataset (like "sarcastic undertone" or "emotionally intense"). These are the kinds of details often overlooked by simpler LLM analyses or manual inspection. For example, DSAI might find that in the news dataset, a certain style of phrasing has a subtle impact on CTR, which wouldn’t be obvious without this kind of analysis. These observations highlight the benefit of a data-centric approach: DSAI can adapt to the particular domain and context of each dataset, rather than relying on one-size-fits-all features. We provide detailed breakdowns and examples from each dataset in Appendix G.
@@
Because DSAI produces human-readable criteria, the extracted features have the potential to readily support various downstream applications such as style transfer (rewriting content to meet certain criteria), generating annotation guidelines (for human labelers to follow), or directly for classification tasks.
To illustrate this, we conducted a toy spam classification experiment using 20 spam and 20 ham samples. Using the five criteria with the highest prominence intensity scores led to effective classification performance, while using the five criteria with the lowest prominence resulted in poor performance (See Figure 3). This small experiment demonstrates that DSAI’s prominence scoring correlates with real utility: the features deemed important by DSAI indeed helped in a classification task, while those deemed unimportant were not useful.